OpenStudy (anonymous):

what is the square root of 14,-3,27,90,-45and 105?

OpenStudy (anonymous):

can u use a calculator?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

sqrt(27) = sqrt(9)xsqrt(3)= 3 sqrt(3)

OpenStudy (anonymous):

sqrt(-3) = 1.73205081 i

OpenStudy (anonymous):

if wanting exact answer then: sqrt(14)=sqrt(14), 3i (imaginary number), 3sqrt(3), 3sqrt(10), (3i)sqrt(5), sqrt(105)=sqrt(105)

OpenStudy (anonymous):

ok thanks and what about 14,27 i am confused and 90 and -45 and 105?

OpenStudy (anonymous):

sqrt(2*7), cannot break down so it stays the same its the exact answer.

OpenStudy (anonymous):

sqrt(3*3*3)=3sqrt(3), remember find the pairs and take them out. only pairs, leave singles inside the sqrt.

OpenStudy (anonymous):

sqrt(3*3*10)=3sqrt(10)

OpenStudy (anonymous):

remember sqrt(-1)=i (imaginary number)

OpenStudy (anonymous):

sqrt(-1*9*5)=(3i)sqrt(5)

OpenStudy (anonymous):

well thanks i will try later

OpenStudy (anonymous):

im 2 confused lol